# InstructAgent: Building User Controllable Recommender via LLM Agent

### v1

@@ 5.1. Recommender System
5.2. Personal Language-based Agent
@@
In the early stages, some researchers (Zhang, 2018; Park et al., 2023; Shanahan et al., 2023) in the NLP field developed dialogue agents with personas to enhance dialogue quality. Language models (Park et al., 2023) are prompted with role descriptions to simulate realistic interactions by storing experiences, synthesizing memories, and dynamically planning actions, resulting in believable individual and social behaviors within interactive environments. WebShop (Yao et al., 2022) attempts to understand product attributes from human-provided text instructions using reinforcement learning and imitation learning. Similar to traditional conversational recommender systems (CRS) (Zhang et al., 2018), it is impractical for users to describe each product attribute every time. With the advancement of large language models (such as GPTs (Achiam et al., 2023)), many researchers (Gur et al., 2023; Deng et al., 2024; Xie et al., 2024; Mei et al., [n. d.]) have begun designing domain-specific agents that integrate various tool learning and memory mechanisms.
@@
More recently, recommendation agents (RecAgent) (Zhao et al., 2024; Wang et al., 2023; Zhang et al., 2024a, b; Wang et al., 2024; Huang et al., 2023a) have been developed to simulate user behaviors and predict user-item interactions. A common design feature among these agents is the use of historical interaction information as user memory (Zhao et al., 2024; Wang et al., 2023; Huang et al., 2023a), with LLMs utilized to generate the ranking results. Unlike platform-side RecAgents, InstructAgent and Instruct2Agent are the first to operate on the user side, generating re-ranking results based on user instructions and individual memory, unaffected by the influence of advantaged users.
APPENDIX
Appendix A Conclusion and Future Direction
In this work, we first establish an instruction-aware recommendation benchmark and design a straightforward instruction-aware agent (InstructAgent) to analyze user instructions and integrate relevant and comprehensive knowledge. Moreover, to enhance the agent’s personalized abilities, we propose individual instruction-aware agent (Instruct2Agent), which incorporates a dynamic memory mechanism to learn from user’s personal feedback and extracts the dynamic interests. In addition to these technical contributions, our work also presents unique and complementary avenues for future research.
More Effective Reranker. In this version of InstructAgent and Instruct2Agent, we construct a zero-shot reranker based on LLMs, such as GPT4-o-mini. Recently, several open-source LLMs (Gunter et al., 2024; Abdin et al., 2024; Team et al., 2024), typically containing fewer model parameters (2-3 billion), have demonstrated strong performance. It is feasible to fine-tune smaller LLMs to build a more effective reranker on our InstructRec dataset. Furthermore, existing advanced recommendation models (Zhai et al., 2024; Xu et al., 2024b) can serve as tools for the agent to retrieve candidate items.
Multi-step Feedback. Although we have constructed various datasets rich in abundant instructions, the feedback for re-ranking results is limited to a single ground-truth item, lacking continuous, multi-step feedback on interactions between users and agents. Additionally, the feedback explanations from users are insufficient. If Instruct2Agent were deployed in a real-world environment, more comprehensive feedback could be collected, enabling the development of more interpretable agents for users.
Mutual Learning. This work builds an agent for users that makes decisions for users and collect feedback from users. The platform-side recommendation models can improve their performance by leveraging the feedback and explanations provided by agents on behalf of their users. Furthermore, recommendation agents (Zhao et al., 2024; Zhang et al., 2024b; Wang et al., 2023; Zhang et al., 2024a) can autonomously and iteratively improve through mutual learning with Instruct2Agent. Moreover, Instruct2Agent can serve as a reward function for RL-based recommendation models (Afsar et al., 2022; Zheng et al., 2018; Wang et al., 2020; Ge et al., 2022b), enhancing their performance.
Appendix B Experiment
B.1. Source Dataset
@@
Amazon Book/Movietv  [^5] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/ (Ni et al., 2019) The Amazon product dataset is a comprehensive repository of consumer reviews and associated metadata, encompassing 142.8 million reviews collected over an 18-year span from May 1996 to July 2014. For our experiments, we leverage two distinct subsets: ”Books” and ”Movies and TV.” Each dataset includes anonymized user and item identifiers, along with user-provided ratings on a 1-5 scale and corresponding textual reviews. Furthermore, rich product metadata is incorporated, such as detailed descriptions, categorical classifications, pricing information, and brand data. This multifaceted dataset provides a fertile ground for both collaborative filtering and content-based recommendation approaches, where the interplay between user behavior, product attributes, and textual feedback can be modeled to advance the state of recommendation systems.
@@
Yelp.  [^7] https://www.kaggle.com/datasets/yelp-dataset/yelp-dataset/versions The Yelp dataset contains over 67,000 reviews focused on businesses, particularly restaurants, from three major English-speaking cities, sourced from the popular Yelp platform. The dataset includes detailed metadata on both businesses and user interactions. Each business is uniquely identified and linked to comprehensive metadata, including its name, geographic location, category (e.g., restaurant, bar, or retail establishment), and additional attributes such as parking availability and reservation policies. This data is invaluable for context-aware recommendation systems, where business features and user feedback intersect to inform personalized recommendations. Anonymized user IDs track user interactions, with additional features such as the number of reviews written, average rating, and social features (e.g., ”friends,” ”useful votes”). Yelp’s textual reviews provide a rich dataset for natural language processing, where the diverse nature of user opinions, combined with structured metadata, offers a robust framework for evaluating and improving context-aware recommendation models.
B.2. Compared Methods
B.2.1. Sequential recommendation methods
@@
B.2.2. Instruction-aware methods
@@ B.2.3. Recommendation Agents
To ensure a fair comparison and optimize computational efficiency, the number of memory-building rounds in AgentCF is set to 1, matching that of our Instruct2Agent. In AgentCF’s experiments, the dataset size is 100, which represents only around 0.1% of the size of our dataset. Moreover, to ensure the generated reranking list without hallucination, we also equipped ToolRec and AgentCF with our self-reflection mechanism.
B.3. Performance Comparison
B.3.1. Echo Chamber Effect
We also report the experimental results evaluating the echo chamber effect in Table 7, Table 8 and Table 9. Ads items are randomly inserted into the candidate ranking list from other domains to simulate advertising scenarios that users may have encountered. To mitigate position bias in LLMs (Liu et al., 2024a), Ads items are added randomly within the candidate list positions. Instruct2Agent accurately identifies users’ instructions and extracts knowledge about their underlying needs, thereby effectively removing undesired Ads. Benefitting from not being trained in a purely data-driven manner and constructing user profiles based on their feedback, our Instruct2Agent also recommends more diverse items to users (both active and less-active items), instead of focusing solely on popular items, and meanwhile improves the overall recommendation performance. Drawing from these experimental results, we conclude that our Instruct2Agent can mitigate the echo chamber effect and act as a protective shield for users.
B.3.2. Protect Less-Active Users
We define the top 20% of users as active, with the remaining 80% classified as less-active (Li et al., 2021; Xu et al., 2023a). Since our data is sampled and filtered using a 10-core process, most users exhibit rich behavioral patterns. Consequently, active users tend to show poorer performance compared to less-active users, largely due to the decline in LLM performance with longer texts (Liu et al., 2024b). As illustrated in Table 10, Table 11 and Table 12, our Instruct2Agent enhances the performance for both active and less-active users. For less-active users, we construct individual profiles based on their feedback, ensuring that these profiles are not influenced by other users. The experimental results demonstrate that our dynamic memory mechanism offers personalized services tailored to each user individually.
Table 7. Evaluation effects (%) of the echo chamber ($\uparrow$) on the InstructRec-Amazon Books. We highlight the methods with the first, second and third best performances.
@@ B.3.2. Protect Less-Active Users
InstructAgent 71.98 67.82 60.74 0.08 38.85 59.51 50.70 57.32
Instruct2Agent 77.15 70.15 64.05 0.09 42.62 64.70 55.25 60.87
Table 8. Evaluation effects (%) of the echo chamber ($\uparrow$) on the InstructRec-Amazon Movietv and InstructRec-GoodReads. We highlight the methods with the first, second and third best performances.
@@
InstructAgent 40.50 60.71 52.11 56.61 23.75 47.50 37.34 42.68
Instruct2Agent 49.51 70.47 61.67 64.69 31.22 57.33 46.23 49.71
Table 9. Evaluation effects (%) of the echo chamber ($\uparrow$) on the InstructRec-Yelp. We highlight the methods with the first, second and third best performances.
@@
InstructAgent 78.24 69.71 56.17 0.12 41.74 62.74 53.82 58.76
Instruct2Agent 87.69 86.20 84.00 0.16 43.67 64.48 55.62 60.20
Table 10. The performance (%) of active and less-active users on InstructRec - Amazon Movietv. We highlight the methods with the first, second and third best performances.
@@
InstructAgent 39.36 57.85 49.98 53.96 34.95 55.19 46.88 51.02
Instruct2Agent 47.32 66.64 58.57 61.22 44.71 64.99 56.60 59.30
Table 11. The performance (%) of active and less-active users on InstructRec - GoodReads. We highlight the methods with the first, second and third best performances.
@@
InstructAgent 24.57 48.12 38.00 43.04 22.62 46.96 36.64 41.70
Instruct2Agent 32.67 58.08 47.28 50.46 29.76 55.39 44.56 48.19
Table 12. The performance (%) of active and less-active users on InstructRec - Yelp. We highlight the methods with the first, second and third best performances.
@@
InstructAgent 37.80 56.17 48.37 52.70 39.40 59.10 50.62 53.90
Instruct2Agent 39.02 58.49 50.23 53.88 43.25 57.75 51.48 56.05
Appendix C Prompt Templates and Examples
@@
C.1. Prompt Templates and Examples Response in InstructAgent
C.1.1. Parser
With the Google Search Tools. The Google Custom Search API operates on a pay-per-use pricing model. The JSON API, used to retrieve web or image search results, charges \$5 per 1,000 queries. There is a limit of 10,000 queries per day.
@@ C.1.1. Parser
C.1.2. Reranker
The reranker uses the following prompt:
@@ C.1.3. Self-reflection
C.2. Prompt Templates and Examples Response in Instruct2Agent
C.2.1. Generator
The generator uses the following prompt:
@@ C.2.1. Generator
C.2.2. Extractor
The extractor uses the following prompt:
@@ C.2.2. Extractor
C.3. Examples of Dataset
C.3.1. Examples in Constructing Dataset
The prompt and examples used in contructing dataset :
@@ C.3.1. Examples in Constructing Dataset
C.3.2. Examples of Filtered Instructions
@@
C.3.3. Examples of Retained Instructions

### v2

@@ B.1 Recommender System
B.2 Personal Language-based Agent
@@
In the early stages, some researchers (Zhang, 2018; Park et al., 2023; Shanahan et al., 2023) in the NLP field developed dialogue agents with personas to enhance dialogue quality. Language models (Park et al., 2023) are prompted with role descriptions to simulate realistic interactions by storing experiences, synthesizing memories, and dynamically planning actions, resulting in believable individual and social behaviors within interactive environments. WebShop (Yao et al., 2022) attempts to understand product attributes from human-provided text instructions using reinforcement learning and imitation learning. Similar to traditional conversational recommender systems (CRS) (Zhang et al., 2018), it is impractical for users to describe each product attribute every time. With the advancement of large language models (such as GPTs (Achiam et al., 2023)), many researchers (Gur et al., 2023; Deng et al., 2024; Xie et al., 2024) have begun designing domain-specific agents that integrate various tool learning and memory mechanisms.
@@
More recently, recommendation agents (RecAgent) (Zhao et al., 2024; Wang et al., 2023; Zhang et al., 2024a, b; Wang et al., 2024; Huang et al., 2023b) have been developed to simulate user behaviors and predict user-item interactions. A common design feature among these agents is the use of historical interaction information as user memory (Zhao et al., 2024; Wang et al., 2023; Huang et al., 2023b), with LLMs utilized to generate the ranking results. Unlike platform-side RecAgents, iAgent and i2Agent are the first to operate on the user side, generating re-ranking results based on user instructions and individual memory, unaffected by the influence of advantaged users.
Appendix C Experiment
C.1 Source Dataset
@@
Amazon Book/Movietv  [^5] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/ (Ni et al., 2019) The Amazon product dataset is a comprehensive repository of consumer reviews and associated metadata, encompassing 142.8 million reviews collected over an 18-year span from May 1996 to July 2014. For our experiments, we leverage two distinct subsets: "Books" and "Movies and TV." Each dataset includes anonymized user and item identifiers, along with user-provided ratings on a 1-5 scale and corresponding textual reviews. Furthermore, rich product metadata is incorporated, such as detailed descriptions, categorical classifications, pricing information, and brand data. This multifaceted dataset provides a fertile ground for both collaborative filtering and content-based recommendation approaches, where the interplay between user behavior, product attributes, and textual feedback can be modeled to advance the state of recommendation systems.
@@
Yelp.  [^7] https://www.kaggle.com/datasets/yelp-dataset/yelp-dataset/versions The Yelp dataset contains over 67,000 reviews focused on businesses, particularly restaurants, from three major English-speaking cities, sourced from the popular Yelp platform. The dataset includes detailed metadata on both businesses and user interactions. Each business is uniquely identified and linked to comprehensive metadata, including its name, geographic location, category (e.g., restaurant, bar, or retail establishment), and additional attributes such as parking availability and reservation policies. This data is invaluable for context-aware recommendation systems, where business features and user feedback intersect to inform personalized recommendations. Anonymized user IDs track user interactions, with additional features such as the number of reviews written, average rating, and social features (e.g., "friends," "useful votes"). Yelp’s textual reviews provide a rich dataset for natural language processing, where the diverse nature of user opinions, combined with structured metadata, offers a robust framework for evaluating and improving context-aware recommendation models.
C.2 Compared Methods
C.2.1 Sequential recommendation methods
@@
C.2.2 Instruction-aware methods
@@ C.2.3 Recommendation Agents
To ensure a fair comparison and optimize computational efficiency, the number of memory-building rounds in AgentCF is set to 1, matching that of our i2Agent. In AgentCF’s experiments, the dataset size is 100, which represents only around 0.1% of the size of our dataset. Moreover, to ensure the generated reranking list without hallucination, we also equipped ToolRec and AgentCF with our self-reflection mechanism.
C.3 Performance Comparison
C.3.1 Echo Chamber Effect
We also report the experimental results evaluating the echo chamber effect in Table 7, Table 8 and Table 9. Ads items are randomly inserted into the candidate ranking list from other domains to simulate advertising scenarios that users may have encountered. To mitigate position bias in LLMs (Liu et al., 2024a), Ads items are added randomly within the candidate list positions. i2Agent accurately identifies users’ instructions and extracts knowledge about their underlying needs, thereby effectively removing undesired Ads. Benefitting from not being trained in a purely data-driven manner and constructing user profiles based on their feedback, our i2Agent also recommends more diverse items to users (both active and less-active items), instead of focusing solely on popular items, and meanwhile improves the overall recommendation performance. Drawing from these experimental results, we conclude that our i2Agent can mitigate the echo chamber effect and act as a protective shield for users.
C.3.2 Protect Less-Active Users
We define the top 20% of users as active, with the remaining 80% classified as less-active (Li et al., 2021; Xu et al., 2023). Since our data is sampled and filtered using a 10-core process, most users exhibit rich behavioral patterns. Consequently, active users tend to show poorer performance compared to less-active users, largely due to the decline in LLM performance with longer texts (Liu et al., 2024b). As illustrated in Table 10, Table 11 and Table 12, our i2Agent enhances the performance for both active and less-active users. For less-active users, we construct individual profiles based on their feedback, ensuring that these profiles are not influenced by other users. The experimental results demonstrate that our dynamic memory mechanism offers personalized services tailored to each user individually.
Table 7: Evaluation effects (%) of the echo chamber ($\uparrow$) on the InstructRec-Amazon Books. We highlight the methods with the first, second and third best performances.
@@ C.3.2 Protect Less-Active Users
iAgent 71.98 67.82 60.74 0.08 38.85 59.51 50.70 57.32
i2Agent 77.15 70.15 64.05 0.09 42.62 64.70 55.25 60.87
Table 8: Evaluation effects (%) of the echo chamber ($\uparrow$) on the InstructRec-Amazon Movietv and InstructRec-GoodReads. We highlight the methods with the first, second and third best performances.
@@
iAgent 40.50 60.71 52.11 56.61 23.75 47.50 37.34 42.68
i2Agent 49.51 70.47 61.67 64.69 31.22 57.33 46.23 49.71
Table 9: Evaluation effects (%) of the echo chamber ($\uparrow$) on the InstructRec-Yelp. We highlight the methods with the first, second and third best performances.
@@
iAgent 78.24 69.71 56.17 0.12 41.74 62.74 53.82 58.76
i2Agent 87.69 86.20 84.00 0.16 43.67 64.48 55.62 60.20
Table 10: The performance (%) of active and less-active users on InstructRec - Amazon Movietv. We highlight the methods with the first, second and third best performances.
@@
iAgent 39.36 57.85 49.98 53.96 34.95 55.19 46.88 51.02
i2Agent 47.32 66.64 58.57 61.22 44.71 64.99 56.60 59.30
Table 11: The performance (%) of active and less-active users on InstructRec - GoodReads. We highlight the methods with the first, second and third best performances.
@@
iAgent 24.57 48.12 38.00 43.04 22.62 46.96 36.64 41.70
i2Agent 32.67 58.08 47.28 50.46 29.76 55.39 44.56 48.19
Table 12: The performance (%) of active and less-active users on InstructRec - Yelp. We highlight the methods with the first, second and third best performances.
@@
iAgent 37.80 56.17 48.37 52.70 39.40 59.10 50.62 53.90
i2Agent 39.02 58.49 50.23 53.88 43.25 57.75 51.48 56.05
Appendix D Prompt Templates and Examples
@@
D.1 Prompt Templates and Examples Response in iAgent
D.1.1 Parser
With the Google Search Tools. [^15] The Google Custom Search API operates on a pay-per-use pricing model. The JSON API, used to retrieve web or image search results, charges \$5 per 1,000 queries. There is a limit of 10,000 queries per day.
@@ D.1.1 Parser
D.1.2 Reranker
@@ D.1.3 Self-reflection
D.2 Prompt Templates and Examples Response in i2Agent
D.2.1 Generator
@@
D.2.2 Extractor
@@
D.3 Examples of Dataset
D.3.1 Examples in Constructing Dataset
@@
D.3.2 Examples of Filtered Instructions
@@
D.3.3 Examples of Retained Instructions
@@
Appendix E Future Directions
E.1 More Effective Reranker.
In this version of iAgent and i2Agent, we construct a zero-shot reranker based on LLMs, such as GPT4-o-mini. Recently, several open-source LLMs (Gunter et al., 2024; Abdin et al., 2024; Team et al., 2024), typically containing fewer model parameters (2-3 billion), have demonstrated strong performance. It is feasible to fine-tune smaller LLMs to build a more effective reranker on our InstructRec dataset. Furthermore, existing advanced recommendation models (Zhai et al., 2024; Xu et al., 2024) can serve as tools for the agent to retrieve candidate items.
E.2 Multi-step Feedback.
Although we have constructed various datasets rich in abundant instructions, the feedback for re-ranking results is limited to a single ground-truth item, lacking continuous, multi-step feedback on interactions between users and agents. Additionally, the feedback explanations from users are insufficient. If i2Agent were deployed in a real-world environment, more comprehensive feedback could be collected, enabling the development of more interpretable agents for users.
E.3 Mutual Learning.
This work builds an agent for users that makes decisions for users and collect feedback from users. The platform-side recommendation models can improve their performance by leveraging the feedback and explanations provided by agents on behalf of their users. Furthermore, recommendation agents (Zhao et al., 2024; Zhang et al., 2024b; Wang et al., 2023; Zhang et al., 2024a) can autonomously and iteratively improve through mutual learning with i2Agent. Moreover, i2Agent can serve as a reward function for RL-based recommendation models (Afsar et al., 2022; Zheng et al., 2018; Wang et al., 2020; Ge et al., 2022b), enhancing their performance.